# New keypoint-based approach for recognising British Sign Language (BSL) from sequences

### v1

@@ 1 Introduction
Our research aims to leverage the power of keypoint representations, which provide valuable insights into the body gestures and expressions used in BSL. By utilizing this approach, we seek to enhance the accessibility and efficiency of BSL recognition, contributing to improved communication for the deaf community. The dataset used in our study encompasses a large-scale collection of diverse BSL expressions, enabling comprehensive training and evaluation of our proposed model. Through this work, we hope to make significant strides in promoting computationally efficient models for British sign language recognition and promoting inclusivity for individuals with hearing impairments.
@@ Acknowledgement
This research was funded by the UKRI’s EPSRC AIMS grant (EP/S024050/1) and we appreciate their support.

### v2

@@ 1 Introduction
Our research aims to leverage the power of keypoint representations, which provide valuable insights into the body gestures and expressions used in BSL. By utilizing this approach, we seek to enhance the accessibility and efficiency of BSL recognition, contributing to improved communication for the deaf community. The dataset used in our study encompasses a large-scale collection of diverse BSL expressions, enabling comprehensive training and evaluation of our proposed model. Through this work, we hope to make significant strides in promoting computationally efficient models for British sign language recognition and fostering inclusivity for individuals with hearing impairments, as leveraging AI technologies can drive enhanced inclusivity across various domains [9], including support for individuals with special needs or disabilities.
@@ Acknowledgement
This research was funded by the UKRI’s EPSRC AIMS grant (EP/S024050/1), and we appreciate their support.